# SPARK: Search Personalization via Agent-Driven Retrieval and Knowledge-sharing

### v1

@@ 4.6. Result Fusion and Answer Synthesis
The fusion stage is coordinated by a central arbiter that merges the per-agent ranked lists into a single aggregated ordering. By default, SPARK employs Reciprocal Rank Fusion (RRF) (cormack2009rrf), a rank aggregation method known for its stability under heterogeneous scoring functions. RRF computes the reciprocal of each document’s rank position (with a tunable offset parameter) across lists and sums them to produce a fused score, ensuring that highly ranked documents from any agent contribute strongly to the final order. Since agents may operate on distinct scoring scales, the arbiter performs confidence calibration (guo2017calibration), mapping raw scores to well-calibrated probability estimates before fusion. Confidence-weighted tie-breaking ensures that no single specialist’s output dominates unfairly, while preserving the benefits of multi-agent diversity. Although RRF provides a strong, domain-agnostic baseline, the system can optionally employ a learned fusion model (karpukhin2020dense; nogueira2019multi) when sufficient labeled data is available, enabling adaptive weighting of agents based on query type and context.
@@
After fusion, a synthesizer persona constructs the final answer. This agent receives the fused top-$k$ documents or passages as structured evidence and generates a coherent, natural-language response grounded in the retrieved content. To ensure transparency, the synthesizer cites source documents inline and flags uncertain claims with hedging markers. The synthesis process incorporates techniques from retrieval-augmented generation (RAG) (lewis2020retrieval) and factuality-aware prompting (shuster2021retrieval) to reduce hallucinations and preserve semantic fidelity to the retrieved evidence.
@@
Personalization can lead to over-narrow retrieval when user profiles dominate ranking preferences. To counteract this, SPARK optimizes for intent coverage during fusion using the Expected Reciprocal Rank – Intent Aware (ERR-IA) metric (chapelle2011intentaware). ERR-IA rewards ranked lists that proportionally represent multiple plausible interpretations of a query, weighted by their estimated likelihood. During agent selection and fusion, this objective encourages inclusion of complementary results, improving resilience on ambiguous or multi-faceted queries. By balancing personalization with diversity, SPARK mitigates the risk of missing relevant content from less dominant but still plausible user intents.
@@ 6. Safety, Privacy, and Governance
Multi-agent, memory-augmented LLM systems such as SPARK introduce safety and governance challenges that surpass those of single-agent models. When multiple agents share a persistent memory store, there is an elevated risk of emergent harmful behavior, cascading reasoning errors, or malicious manipulation of shared state. An agent’s long-term memory can be poisoned by adversarial inputs, leading to inaccurate outputs, or probed via targeted prompt injections designed to elicit sensitive details (carlini2023extracting; shi2023prompt). Recent work demonstrates that LLM agents are particularly vulnerable to memory extraction attacks, where carefully crafted conversational queries progressively siphon stored personal data from long-term memory (wang2025memoryrisk). To mitigate such risks, data minimization is critical—SPARK should retain only essential information, for the shortest necessary duration, consistent with privacy-by-design principles. Techniques such as user-led pruning or anonymization, as implemented in Rescriber (zhou2025rescriber), enable end-users to redact personally identifiable information before it is committed to memory. Beyond deletion, systems can apply pseudonymization, encryption-at-rest, and contextual hashing to reduce data sensitivity. Coupled with transparent policy disclosures, this ensures user control over what persists in SPARK’s semantic memory. Furthermore, memory retrieval operations should be fully auditable, with provenance tracking for each stored item and an explainable record of why a given memory item influenced current agent outputs. Such explainability supports AI governance requirements, improves accountability, and enables targeted remediation in the event of malicious manipulation.
@@
Preventing sensitive information leakage in multi-agent LLM systems requires a combination of technical safeguards and procedural governance. On the technical side, all inter-agent communication should be scope-limited: agents should only exchange abstracted insights or embeddings, never raw personal data. Sensitive entries in long-term memory can be selectively encrypted, hashed, or replaced with semantically equivalent but anonymized forms prior to storage (yuan2024privacyrag). Research on retrieval-augmented LLMs has shown that without such safeguards, stored context can leak private details and enable attacks such as membership inference (jiang2024ragthief). To guard against these risks, strict role-based access control (RBAC) can be applied, ensuring only authorized agents retrieve certain classes of data. Procedurally, organizations deploying SPARK should implement adversarial audits (sandvig2014audit), where red-team testers probe the system with crafted inputs to detect both inadvertent and intentional data leaks. These audits, combined with automated privacy scanning pipelines, can detect whether agents are inadvertently exposing user data in outputs. Governance policies should require regular memory purges, retention-limit enforcement, and clear escalation processes for incidents. A compliance log—recording retrieval events, agent decisions, and data flows—supports forensic investigation and aligns with emerging AI governance frameworks.
@@
Another significant governance challenge is personalization harm, where SPARK’s semantic memory over-optimizes for a user’s established profile, narrowing information exposure. This filter bubble effect has been extensively studied in recommender systems (nguyen2014filterbubble; pariser2011filterbubble), where algorithmic reinforcement leads to a progressive narrowing of presented content. Over time, such narrowing can reduce diversity, entrench bias, and undermine user autonomy. To counteract this, SPARK should integrate diversity-promoting objectives into retrieval and ranking. For example, incorporating the Expected Reciprocal Rank for Intent-Awareness (ERR-IA) (chapelle2011erria) during result fusion ensures that top-ranked outputs cover multiple plausible interpretations of the query, proportionate to their estimated likelihood. Moreover, fairness-aware memory update policies can prevent over-weighting transient or repetitive experiences. Techniques such as controlled forgetting, counterfactual augmentation, and entropy-based diversity constraints help maintain a balanced profile representation. SPARK could also introduce serendipity injections—intentionally surfacing out-of-profile but potentially relevant information—to mitigate over-personalization. Finally, long-term evaluation of personalization strategies should consider not only click-through rates but also user satisfaction, content diversity, and exposure to novel perspectives (smith2022fairpersonalization). In summary, safety, privacy, and governance in SPARK require a multi-pronged approach: strict data minimization, secure and auditable memory access, procedural oversight, and bias-aware personalization safeguards.
@@ 7. Limitations and Risks
While SPARK’s multi-agent architecture enables specialization, decomposition, and collaborative reasoning, it can introduce unnecessary complexity for low-difficulty tasks. Coordinating multiple LLM-based agents—such as a planner, retriever, and summarizer—requires additional message passing, synchronization, and aggregation. For straightforward queries, this often results in longer latency and higher computational cost compared to a single, well-tuned agent (martyr2025coordination). Each agent invocation adds potential failure points: message truncation, misinterpretation of intermediate results, or deadlocks in agent pipelines. Empirical benchmarks suggest that as coordination depth increases, diminishing returns can set in, with performance parity or even degradation relative to single-agent baselines on bounded tasks (mirchandani2024dawn). Frameworks like DAWN propose gating mechanisms to dynamically bypass multi-agent orchestration when predicted task complexity is low, enabling a fallback to single-agent or even non-LLM deterministic modules for optimal efficiency. This adaptivity reduces overhead while preserving multi-agent benefits for truly complex problems, balancing throughput with reasoning capability.
@@
SPARK’s semantic memory continuously adapts based on user interactions, but without safeguards, it risks personalization drift—where transient user behavior is mistaken for stable preference (dou2007session). For example, a user researching an unrelated topic for a short-term task may inadvertently skew their long-term profile if the system overweights recent clicks or queries. Over time, this drift can degrade relevance, reinforce biases, and create self-perpetuating filter bubbles (nguyen2014filterbubble; pariser2011filterbubble). To mitigate drift, SPARK can apply decay-based weighting, requiring repeated evidence before committing profile changes, and separating session-level signals from long-term memory updates. Diversity-aware retrieval objectives, such as ERR-IA (chapelle2011erria), can further ensure the result set remains varied even when the profile is narrowly defined. Periodic model audits—either algorithmic or with human oversight—can detect and correct misaligned profiles. Additionally, fairness-aware personalization strategies (smith2022fairpersonalization) help prevent overrepresentation of historically dominant content, supporting a more balanced and exploratory user experience.
@@
SPARK’s layered memory design inherently stores potentially sensitive data, including past queries, retrieved documents, and inferred user attributes. Without strong controls, these can become a target for adversarial exploitation. Memory extraction and retrieval-augmented leakage attacks, such as RAG-Thief (jiang2024ragthief), demonstrate that crafted prompts can induce models to reveal stored private data. Mitigations include encryption-at-rest, in-memory sanitization pipelines, and scope-restricted retrieval—where agents only access the minimum information necessary for their task (yuan2024privacyrag). Differential privacy techniques can further limit exposure by perturbing stored representations while preserving utility for personalization. Procedurally, a privacy governance layer should enforce retention limits, provide user-facing dashboards for history inspection and deletion, and log all memory access for auditability. Similar to practices in safety-critical AI, adversarial red-teaming (sandvig2014audit) should be regularly conducted to probe for unintended disclosure channels. Abstracting stored data into vector embeddings rather than raw text reduces identifiability while maintaining retrieval performance, striking a balance between utility and confidentiality.
@@
Evaluating multi-agent LLM systems like SPARK requires more than traditional accuracy metrics; it must also account for coordination efficiency, fairness, robustness, and trustworthiness over time (safeagentbench2024). Complex systems face knowledge divergence between the base model’s static parameters and the evolving user-specific memory (peng2023unlearning), leading to potential contradictions in outputs. Continuous learning pipelines should support unlearning—removing or correcting erroneous memory entries—while maintaining consistency with the base model’s knowledge. Longitudinal evaluation can track personalization benefits alongside diversity preservation, bias mitigation, and privacy outcomes. Finally, maintenance of such systems demands active monitoring of agent performance, with automatic fallback strategies when degraded behavior is detected. By embedding evaluation and maintenance into the operational lifecycle, SPARK can sustain performance while minimizing the risks inherent to complex, memory-augmented multi-agent systems.

### v2

@@ 4.6. Result Fusion and Answer Synthesis
The fusion stage is coordinated by a central arbiter that merges the per-agent ranked lists into a single aggregated ordering. By default, SPARK employs Reciprocal Rank Fusion (RRF) (Cormack et al., 2009), a rank aggregation method known for its stability under heterogeneous scoring functions. RRF computes the reciprocal of each document’s rank position (with a tunable offset parameter) across lists and sums them to produce a fused score, ensuring that highly ranked documents from any agent contribute strongly to the final order. Since agents may operate on distinct scoring scales, the arbiter performs confidence calibration (Guo et al., 2017), mapping raw scores to well-calibrated probability estimates before fusion. Confidence-weighted tie-breaking ensures that no single specialist’s output dominates unfairly, while preserving the benefits of multi-agent diversity. Although RRF provides a strong, domain-agnostic baseline, the system can optionally employ a learned fusion model (Karpukhin et al., 2020; Nogueira and Cho, 2019) when sufficient labeled data is available, enabling adaptive weighting of agents based on query type and context.
@@
After fusion, a synthesizer persona constructs the final answer. This agent receives the fused top-$k$ documents or passages as structured evidence and generates a coherent, natural-language response grounded in the retrieved content. To ensure transparency, the synthesizer cites source documents inline and flags uncertain claims with hedging markers. The synthesis process incorporates techniques from retrieval-augmented generation (RAG) (Lewis et al., 2020) and factuality-aware prompting (Shuster et al., 2021) to reduce hallucinations and preserve semantic fidelity to the retrieved evidence.
@@
Personalization can lead to over-narrow retrieval when user profiles dominate ranking preferences. To counteract this, SPARK optimizes for intent coverage during fusion using the Expected Reciprocal Rank – Intent Aware (ERR-IA) metric (Chapelle and Zhang, 2011). ERR-IA rewards ranked lists that proportionally represent multiple plausible interpretations of a query, weighted by their estimated likelihood. During agent selection and fusion, this objective encourages inclusion of complementary results, improving resilience on ambiguous or multi-faceted queries. By balancing personalization with diversity, SPARK mitigates the risk of missing relevant content from less dominant but still plausible user intents.
@@ 6. Safety, Privacy, and Governance
Multi-agent, memory-augmented LLM systems such as SPARK introduce safety and governance challenges that surpass those of single-agent models. When multiple agents share a persistent memory store, there is an elevated risk of emergent harmful behavior, cascading reasoning errors, or malicious manipulation of shared state. An agent’s long-term memory can be poisoned by adversarial inputs, leading to inaccurate outputs, or probed via targeted prompt injections designed to elicit sensitive details (Carlini and others, 2023; Shi and et al., 2023). Recent work demonstrates that LLM agents are particularly vulnerable to memory extraction attacks, where carefully crafted conversational queries progressively siphon stored personal data from long-term memory (Wang et al., 2025). To mitigate such risks, data minimization is critical—SPARK should retain only essential information, for the shortest necessary duration, consistent with privacy-by-design principles. Techniques such as user-led pruning or anonymization, as implemented in Rescriber (Zhou et al., 2025), enable end-users to redact personally identifiable information before it is committed to memory. Beyond deletion, systems can apply pseudonymization, encryption-at-rest, and contextual hashing to reduce data sensitivity. Coupled with transparent policy disclosures, this ensures user control over what persists in SPARK’s semantic memory. Furthermore, memory retrieval operations should be fully auditable, with provenance tracking for each stored item and an explainable record of why a given memory item influenced current agent outputs. Such explainability supports AI governance requirements, improves accountability, and enables targeted remediation in the event of malicious manipulation.
@@
Preventing sensitive information leakage in multi-agent LLM systems requires a combination of technical safeguards and procedural governance. On the technical side, all inter-agent communication should be scope-limited: agents should only exchange abstracted insights or embeddings, never raw personal data. Sensitive entries in long-term memory can be selectively encrypted, hashed, or replaced with semantically equivalent but anonymized forms prior to storage (Yuan et al., 2024). Research on retrieval-augmented LLMs has shown that without such safeguards, stored context can leak private details and enable attacks such as membership inference (Jiang et al., 2024). To guard against these risks, strict role-based access control (RBAC) can be applied, ensuring only authorized agents retrieve certain classes of data. Procedurally, organizations deploying SPARK should implement adversarial audits (Sandvig et al., 2014), where red-team testers probe the system with crafted inputs to detect both inadvertent and intentional data leaks. These audits, combined with automated privacy scanning pipelines, can detect whether agents are inadvertently exposing user data in outputs. Governance policies should require regular memory purges, retention-limit enforcement, and clear escalation processes for incidents. A compliance log—recording retrieval events, agent decisions, and data flows—supports forensic investigation and aligns with emerging AI governance frameworks.
@@
Another significant governance challenge is personalization harm, where SPARK’s semantic memory over-optimizes for a user’s established profile, narrowing information exposure. This filter bubble effect has been extensively studied in recommender systems (Nguyen et al., 2014; Pariser, 2011), where algorithmic reinforcement leads to a progressive narrowing of presented content. Over time, such narrowing can reduce diversity, entrench bias, and undermine user autonomy. To counteract this, SPARK should integrate diversity-promoting objectives into retrieval and ranking. For example, incorporating the Expected Reciprocal Rank for Intent-Awareness (ERR-IA) (Chapelle et al., 2011) during result fusion ensures that top-ranked outputs cover multiple plausible interpretations of the query, proportionate to their estimated likelihood. Moreover, fairness-aware memory update policies can prevent over-weighting transient or repetitive experiences. Techniques such as controlled forgetting, counterfactual augmentation, and entropy-based diversity constraints help maintain a balanced profile representation. SPARK could also introduce serendipity injections—intentionally surfacing out-of-profile but potentially relevant information—to mitigate over-personalization. Finally, long-term evaluation of personalization strategies should consider not only click-through rates but also user satisfaction, content diversity, and exposure to novel perspectives (Smith et al., 2022). In summary, safety, privacy, and governance in SPARK require a multi-pronged approach: strict data minimization, secure and auditable memory access, procedural oversight, and bias-aware personalization safeguards.
@@ 7. Limitations and Risks
While SPARK’s multi-agent architecture enables specialization, decomposition, and collaborative reasoning, it can introduce unnecessary complexity for low-difficulty tasks. Coordinating multiple LLM-based agents—such as a planner, retriever, and summarizer—requires additional message passing, synchronization, and aggregation. For straightforward queries, this often results in longer latency and higher computational cost compared to a single, well-tuned agent (Martyr, 2025). Each agent invocation adds potential failure points: message truncation, misinterpretation of intermediate results, or deadlocks in agent pipelines. Empirical benchmarks suggest that as coordination depth increases, diminishing returns can set in, with performance parity or even degradation relative to single-agent baselines on bounded tasks (Mirchandani et al., 2024). Frameworks like DAWN propose gating mechanisms to dynamically bypass multi-agent orchestration when predicted task complexity is low, enabling a fallback to single-agent or even non-LLM deterministic modules for optimal efficiency. This adaptivity reduces overhead while preserving multi-agent benefits for truly complex problems, balancing throughput with reasoning capability.
@@
SPARK’s semantic memory continuously adapts based on user interactions, but without safeguards, it risks personalization drift—where transient user behavior is mistaken for stable preference (Dou et al., 2007b). For example, a user researching an unrelated topic for a short-term task may inadvertently skew their long-term profile if the system overweights recent clicks or queries. Over time, this drift can degrade relevance, reinforce biases, and create self-perpetuating filter bubbles (Nguyen et al., 2014; Pariser, 2011). To mitigate drift, SPARK can apply decay-based weighting, requiring repeated evidence before committing profile changes, and separating session-level signals from long-term memory updates. Diversity-aware retrieval objectives, such as ERR-IA (Chapelle et al., 2011), can further ensure the result set remains varied even when the profile is narrowly defined. Periodic model audits—either algorithmic or with human oversight—can detect and correct misaligned profiles. Additionally, fairness-aware personalization strategies (Smith et al., 2022) help prevent overrepresentation of historically dominant content, supporting a more balanced and exploratory user experience.
@@
SPARK’s layered memory design inherently stores potentially sensitive data, including past queries, retrieved documents, and inferred user attributes. Without strong controls, these can become a target for adversarial exploitation. Memory extraction and retrieval-augmented leakage attacks, such as RAG-Thief (Jiang et al., 2024), demonstrate that crafted prompts can induce models to reveal stored private data. Mitigations include encryption-at-rest, in-memory sanitization pipelines, and scope-restricted retrieval—where agents only access the minimum information necessary for their task (Yuan et al., 2024). Differential privacy techniques can further limit exposure by perturbing stored representations while preserving utility for personalization. Procedurally, a privacy governance layer should enforce retention limits, provide user-facing dashboards for history inspection and deletion, and log all memory access for auditability. Similar to practices in safety-critical AI, adversarial red-teaming (Sandvig et al., 2014) should be regularly conducted to probe for unintended disclosure channels. Abstracting stored data into vector embeddings rather than raw text reduces identifiability while maintaining retrieval performance, striking a balance between utility and confidentiality.
@@
Evaluating multi-agent LLM systems like SPARK requires more than traditional accuracy metrics; it must also account for coordination efficiency, fairness, robustness, and trustworthiness over time (Li et al., 2024). Complex systems face knowledge divergence between the base model’s static parameters and the evolving user-specific memory (Peng et al., 2023), leading to potential contradictions in outputs. Continuous learning pipelines should support unlearning—removing or correcting erroneous memory entries—while maintaining consistency with the base model’s knowledge. Longitudinal evaluation can track personalization benefits alongside diversity preservation, bias mitigation, and privacy outcomes. Finally, maintenance of such systems demands active monitoring of agent performance, with automatic fallback strategies when degraded behavior is detected. By embedding evaluation and maintenance into the operational lifecycle, SPARK can sustain performance while minimizing the risks inherent to complex, memory-augmented multi-agent systems.